OpenStudy (anonymous):

If y^5 +3x^2y^2 +5x^4 = 49, then dy/dx at the point (-1, 2) is?

OpenStudy (anonymous):

ok so you want to use implicit differentiation differentiate both sides dy/dx[y^5] = ?

OpenStudy (anonymous):

you can't differentiate normally. is dy/dx[y^5] really = 5y^4? of course it's not, that wouldn't make sense. we use the chain rule. take the derivative of y^5 with respect to y then multiply it by dy/dx to get the derivative of dy/dx[y^5], ya feel

OpenStudy (anonymous):

Do we set the equation equal to zero first? so that 45 isnt on one side

OpenStudy (anonymous):

we just differentiate both sides, so eventually the right hand side will be 0 bc d/dx[constant] = 0

OpenStudy (anonymous):

I am not sure :/

OpenStudy (anonymous):

@urb0b (20x^3 + 6xy^2)/(6x^2y+5y^4)

OpenStudy (anonymous):

to do implicit differentiation, you do stuff like this: